OpenStudy (anonymous):

6m^3-16m^2+15m-40/2m^2+5

OpenStudy (anonymous):

A. 3m - 8 B. 3m + 5 C. 2m - 6 D. 2m + 8

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

6m^3-16m^2+15m-40 (6m^3-16m^2)+(15m-40) 2m^2(3m-8)+(15m-40) I'll let you finish factoring

jimthompson5910 (jim_thompson5910):

after you factor, you'll find cancellations occur and the answer will be left behind

OpenStudy (anonymous):

would it be A?

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

Thanks

jimthompson5910 (jim_thompson5910):

sure thing

OpenStudy (anonymous):

$$\begin{align*}\frac{6m^3-16m^2+15m-40}{2m^2+5}&=\frac{2m^2(3m-8)+5(3m-8)}{2m^2+5}\\&=\frac{(3m-8)(2m^2+5)}{2m^2+5}\\&=3m-8\end{align*}$$